A chemist has prepared two acid solutions, one of which is 4% acid by volume, another 11% acid. How many cubic centimeters of each should the chemist mix together to obtain 50cm3 of 6.24% solution?
Answer by checkley77(12844)

.11x+.04(50-x)=.0624*50
.11x+2-.04x=3.12
.07x=3.12-2
.07x=1.12
x=1.12/.07
x=16 cm of 11% solution.
50-16=34 cm of 4% solution.
Proof:
.11*16+.04*34=3.12
1.76+1.36=3.12
3.12=3.12